AaBbc clodH AaBbCCD Heading 1 Cobalt Chloride Sodium Fluoride ID Chromate Monohydrate 3. In an experiment like the one you performed in lab, a hydrated Cuso4 salt was heated and thus dehydrated. Determine the H20 by weight in the hydrated CuSO4 salt. (a) mass of empty dish 56.123 g (b) mass of dish t hydrated salt 56.978 g (c) mass of disht dehydrated salt-56.719 g Work Answer: 56.978 g. 56.123 g. 0.855 g, hydrate 30.292% water 56.719 g. 56,123 g. 0.596 g. dehydrate 0.855 g. -0.596 g 0.259 g. water loss 0.259 g. 0.855 g. x 100 30.292% water 4. Use the data found in question number 3: A. Determine the empirical formula for the hydrated Cusoe salt. Work: rical Formula B State the chemical name of this hvdrated salt below
Answer #1

A . According to data which you have provided we can easily calculate empirical formula of hydrated salt

if percentage of water is 30.292

then percentage of CuSO4 = 100 - 30.292 = 69.708

now number of moles of water = 30.292/18 = 1.68228

number of moles of CuSO4 = 69.708 / 159.5 = 0.43704

now you can find out the number of moles of both the components in whole number by dividing with the least digit number

ratio of number of moles is CuSO4​ H2O

0.43704/0.43704 1.68228/0.43704

= 1 =3.8

It means the ratio is approx 1: 4 , so the molecular formula of hydrated salt should be CuSO4.4H2O

NOTE      Molecular formula of hydrated Copper sulphate is CuSO4.5H2O. there is no any mistake in calculation , I think there is some problem in your data, please check it and solve the question by above method .

B. Chemical name of this compound is Copper Sulphate ( hydrated)